OpenStudy (hpfan101):

Evaluate the difference quotient for the given function. Simplify your answer. f(x)=1/x, f(x)- f(a)/x-a Can someone please show me the steps? I know the answer but I don't quite get how to solve it. I tried but kept getting the wrong simplified answer.

OpenStudy (anonymous):

f(x) = 1/x Then ((f(x) - f(a))/(x-a) = (1/x - 1/a)/(x-a) Multiply everything by ax/ax we have (a - x)/(ax(x-a)) = -(a-x)/(ax*(a-x)) = -1/(ax) which is the answer

OpenStudy (hpfan101):

Oh, I get it now. Thank you!! :)